Sharon has 30 nickels and dimes having a total value of $2.35. How many of each kind does she have?
Answer by jorel1380(3719)

n+d=30
5n+10d=235
5n+5d=150
5d=85
d=17
n=13
Sharon has 13 nickels and 17 dimes..